OpenStudy (anonymous):

How do I solve k+1=3k-1

OpenStudy (anonymous):

3k-1=k+1 you then subtract k from k+1 you then get this equation 2k-1=1 you then add 1 so you get 2k by itself you then get this equation 2k=2 then divide by 2 2k/2=2/2 so k=1

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

No problem1

OpenStudy (anonymous):

Can you help me with some more please ?

OpenStudy (anonymous):

sure